Find the GCF of the product. (3b + 6)(2b – 10)
I came up with 5 Is it right?

Answer by funmath(2933)

Find the GCF of the product. (3b + 6)(2b – 10)
You can factor out a 3 from the 1st set of parentheses: 3(b+2)
and a 2 from the 2nd set of parentheses: 2(b-5)
Factors are multiplied, not added: 3(b+2)*2(b-5)=6(b+2)(b-5)
The GCF is 6.
:
Proof:
(3b+6)(2b-10)=6b^2-30b+12b-60=6b^2-18b-60
The GCF is 6: 6(b^2-3b-10)
If you factor it further you get: 6(b-5)(b+2)
Happy Calculating!!!